# Linear Bandits with Partially Observable Features

### v2

@@ 4.2 Doubly Robust Lasso Estimator
where $\check{\boldsymbol{\mu}}_{t}^{L}$ is the imputation estimator that fills in the missing rewards of unselected arms in round $t$, as defined in Eq. 8.
For $a\neq\widetilde{a}_{t}$, i.e., an arm that is not selected in the round $t$, we impute the missing rewards using $\widetilde{\mathbf{x}}_{k}^{\top}\check{\boldsymbol{\mu}}_{t}^{L}$. For $a=\widetilde{a}_{t}$, on the other hand, the term $\mathbb{I}(\widetilde{a}_{t}=a)y_{a,t}/\phi_{a,t}$ calibrates the predicted reward to ensure the unbiasedness of the pseudo-rewards for all arms. Given that $\mathbb{E}_{\widetilde{a}_{t}}[\mathbb{I}(\widetilde{a}_{t}=a)]=\mathbb{P}(\widetilde{a}_{t}=a)=\phi_{a,t}$, taking the expectation over $\widetilde{a}_{t}$ on both sides of Eq. 10 gives $\mathbb{E}_{\widetilde{a}_{t}}[\widetilde{y}_{a,t}]=\mathbb{E}_{t-1}[y_{a,t}]=\widetilde{\mathbf{x}}_{a}^{\top}{{\boldsymbol{\mu}}_{\star}}$ for all $a\in[K]$. Although the estimate $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$ may have a large error, it is multiplied by the mean-zero random variable $(1-\mathbb{I}(\widetilde{a}_{t}=a)/\phi_{a,t})$, which makes the resulting pseudo-rewards defined in Eq. 10 robust to errors of $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$. The pseudo-rewards can only be computed—and thus can only be used—when the pseudo-action $\widetilde{a}_{t}$ matches the actual action $a_{t}$, that is, when the event $\mathcal{M}_{t}$ occurs. Since $\mathcal{M}_{t}$ occurs with high probability, we are able to compute pseudo-rewards for almost all rounds.
@@
and the following theorem provides a theoretical guarantee that this estimator converges to the across all arms after a sufficient number of exploration rounds.
@@
Let $d_{h}$ denote the dimension of the projected latent rewards defined in Eq. 6. Then for all round $t$ such that $t\geq|\mathcal{E}_{t}|$, with probability at least $1-2\delta/t^{2}$,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star})|\leq\frac{8\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}}.$
Although the DR Lasso estimator leverages $K$-dimensional feature vectors, its error bound depends only logarithmically on $K$. Such fast convergence is typically achieved under classical regularity conditions, including the compatibility condition and the restrictive minimum eigenvalue condition (van de Geer & Bühlmann, 2009; Bühlmann & van de Geer, 2011). Existing Lasso-based bandit approaches (Kim & Paik, 2019; Bastani & Bayati, 2020; Oh et al., 2021; Ariu et al., 2022; Chakraborty et al., 2023; Lee et al., 2025) generally impose these conditions directly on the feature vectors. In contrast, our approach does not require such assumptions, as the augmented features are orthogonal vectors lying in $\mathrm{R}(\mathbf{X})^{\perp}$. Moreover, their average Gram matrix satisfies $\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\geq\min\{1,\lambda_{\min}(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top})\}$. Thus, the convergence rate scales only as $\sqrt{\log K}$ with respect to $K$.
The consistency is proved by bounding the two components of the error in the pseudo-rewards defined in Eq. 10: (i) the noise of the reward and (ii) the error of the imputation estimator $\check{\boldsymbol{\mu}}_{t}$. Since (i) is sub-Gaussian, it can be bounded using martingale concentration inequalities. For (ii), the imputation error $\widetilde{\mathbf{x}}_{a}^{\top}(\check{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star})$ is multiplied by the mean-zero random variable $\left(1-{\mathbb{I}(\widetilde{a}_{t}=a)}/{\phi_{a,t}}\right)$ and thus the magnitude of the error can be bounded by $\|\check{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star}\|_{1}/\sqrt{t}$. The complete proof is deferred to Section E.2.
@@
12:        Sample $\widetilde{a}_{t}$ according to  Eq. 9.
@@
19:        Update $\widehat{\boldsymbol{\mu}}^{L}_{t}$ following Eq. 11 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}^{L}_{t}$ following Eq. 8.
@@ 5.1 Robust to Latent Features (RoLF) Algorithm
In this section, we present our algorithm RoLF in Algorithm 1. In the initialization step, given the observed features, our algorithm constructs a set of orthogonal basis vectors $\{\mathbf{b}_{1}^{\top},\dots,\mathbf{b}_{K-d}^{\top}\}\subseteq\mathrm{R}(\mathbf{X})^{\perp}$ to augment each observed feature vector. The algorithm first selects a candidate action $\widehat{a}_{t}$—either uniformly at random or based on the estimated reward—and then resamples $a_{t}$ and $\widetilde{a}_{t}$ until they match or the maximum number of trials allowed for each round, $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/p)$, is reached. Once the resampling phase ends, the agent selects $a_{t}$ and the corresponding reward $y_{a_{t},t}$ is observed. If $\widetilde{a}_{t}$ and $a_{t}$ match within $\rho_{t}$, then both the imputation and the main estimators are updated; otherwise, neither is updated.
Our proposed algorithm does not require knowledge of the dimension of the unobserved features $d_{u}$, nor the dimension of the reward component projected onto the orthogonal complement of $\mathrm{R}(\mathbf{X})^{\perp}$. Although we present the algorithm for fixed feature vectors, the algorithm is also applicable to time-varying feature vectors. In such cases, we estimate the bias caused by unobserved features by augmenting the standard basis. For further details, refer to Appendix D.
@@ 5.2 Regret Analysis
Let $d_{h}$ denote the number of nonzero coefficients in the representation of the projected latent reward as defined in Eq. 6. Then for any $\delta\in(0,1)$ and $p\in(1/2,1)$, with probability at least $1-6\delta$, the cumulative regret of the proposed algorithm is bounded by
$\begin{split}\mathrm{Reg}(T)&\leq 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}\\ &\quad+\frac{4\sqrt{T}}{K-1}\frac{\log((T+1)^{2}/\delta)}{\log(1/p)}+4\sqrt{2T\log\frac{2}{\delta}}\\ &\quad+4\delta+\frac{16\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}}.\end{split}$
To the best of our knowledge, Theorem 3 provides the first regret bound that converges faster than $\widetilde{O}(\sqrt{KT})$, specifically for algorithms that account for unobserved features without relying on any structural assumptions. Assuming $\|\widetilde{\mathbf{x}}_{a}\|_{\infty}\leq 1$ (instead of $\|\widetilde{\mathbf{x}}_{a}\|_{2}\leq 1$), $\widetilde{\sigma}^{2}_{\min}$ and $\widetilde{\sigma}_{\max}^{2}$ are constant independent of $d$ or $K$. Note that the number of rounds required for the exploration phase is $O(K^{3}\log KT)$, which grows only logarithmic with the time horizon $T$. The factor $K^{3}$ is irreducible, as the algorithm must estimate all $K$ biases from the missing features. By employing the Gram matrix of the augmented feature vectors over the action space $\mathcal{A}$, $\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$, in combination with DR estimation, we reduce the required exploration phase time from $O(K^{4}\log KT)$ to $O(K^{3}\log KT)$, thereby improving the sample complexity by a factor of $K$.
As demonstrated in Theorem 2, the last term on the right-hand side of the regret bound is proportional to $\sqrt{d+d_{h}}$ rather than $\sqrt{K}$, resulting in an overall regret bound of $O(\sqrt{(d+d_{h})T\log KT})$. The specific value of $d_{h}$ depends on the relationship between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$, as discussed in Section 4.1. Table 2 summarizes how the regret bound varies under different relationships between these subspaces. The formal proof of Theorem 3 is deferred to Section E.3.
@@ 6.1 Experimental Setup
In this experiment, we simulate and compare our algorithms, Algorithm 1 and Algorithm 2 (Appendix C), with baseline OLB algorithms: LinUCB (Li et al., 2010) and LinTS (Agrawal & Goyal, 2013). These algorithms adopt UCB and Thompson sampling strategies, respectively, assuming a linear reward model based on observed features. We also include DRLasso (Kim & Paik, 2019) since our algorithm employs DR estimation with Lasso estimator, and UCB($\delta$) (Lattimore & Szepesvári, 2020), an MAB algorithm that ignores features, to assess the effectiveness of using feature information, even under partial observability.
@@
In the simulation, after construction of the true features $\mathbf{z}_{a}\in\mathbb{R}^{d_{z}}$ and observed features $\mathbf{x}_{a}\in\mathbb{R}^{d}$, we apply singular value decomposition (SVD) to the observed feature matrix $\mathbf{X}$ to derive orthogonal basis vectors $\{\mathbf{b}_{1}^{\top},\dots,\mathbf{b}_{K-d}^{\top}\}$ orthogonal to $\mathrm{R}(\mathbf{X})$. These vectors are then concatenated with $\mathbf{X}$ to form the augmented feature matrix. The reward parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$ is sampled from $\mathrm{Unif}(-1/2,1/2)$, and the rewards are generated following Eq. 2. For the hyperparamters in our algorithms, the coupling probability $p$ and the confidence parameter $\delta$, are set to $0.6$ and $10^{-4}$, respectively. The total decision horizon is $T=1200$. Throughout the experiments, we fix the number of arms at $K=30$ and the dimension of the true features at $d_{z}=35$, ensuring $d_{z}\geq d$ to cover both scenarios. Further details on the experimental setup are provided in Appendix B.
@@ 6.2.1 Scenario 1
In this scenario, $d$ is set to $\lfloor d_{z}/2\rfloor$, so that the agent observes only about half of the full feature dimension. From Figure 2, we observe that the baseline OLB algorithms—LinUCB, LinTS, and DRLasso—perform worse than our algorithms in terms of both the level and robustness of cumulative regret. This tendency is consistently observed across all three cases, implying that linear bandit algorithms fail to identify the optimal arm when relying solely on observed features in environments with unobserved components, since the algorithms cannot capture the portion of the reward attributed to the unobserved features. However, our algorithms show almost the same performance regardless of the relationship between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$, indicating the robustness to variations in feature observability structure.
Note that for all cases our algorithms—RoLF-Lasso (Algorithm 1) and RoLF-Ridge (Algorithm 2)—exhibit a sharp decline in the growth rate of cumulative regret after a certain number of rounds. This behavior is primarily due to the forced-exploration phase built into the algorithms, which ensures sufficient coverage of the action space in the early stages. Following this phase, the resampling and coupling strategies further enhance the efficiency of the DR estimation, leading to slower regret accumulation over time. This pattern is commonly shown from experimental results of other bandit algorithms employing the forced-exploration strategy (Goldenshluger & Zeevi, 2013; Hao et al., 2020; Chakraborty et al., 2023).
Furthermore, in Case 2 (Figure 2(b)), the cumulative regret grows more slowly than in other cases. This behavior is explained by the relationship $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$, implying the reward components projected onto $\mathrm{R}(\mathbf{X})^{\perp}$ can be fully expressed by the observed features. Hence, the augmented features fully capture the underlying reward structure, enabling faster convergence relative to the other scenarios.

### v3

@@ 4.2 Doubly Robust Lasso Estimator
where $\check{\boldsymbol{\mu}}_{t}^{L}$ is the imputation estimator that fills in the missing rewards of unselected arms at round $t$, as defined in Eq.˜8.
As shown in Eq.˜10, the pseudo-reward includes an inverse probability term. Hence, when DR estimation is applied under the $\epsilon_{t}$-greedy policy with $\epsilon_{t}=t^{-1/2}$, its variance can grow unbounded over time. To mitigate this issue, we propose a resampling and coupling strategy: by coupling the $\epsilon_{t}$-greedy policy with the multinomial distribution (Eq.˜9), we ensure that each inverse probability weight $\phi_{a,t}^{-1}$, for $a\in[K]$, remains bounded by $O(K)$. This strategy effectively stabilizes the variance of the pseudo-rewards.
Moreover, one can show that the resampling succeeds with high probability for each round. Let $\mathcal{M}_{t}$ denote the event that the pseudo-action $\widetilde{a}_{t}$ matches the chosen action $a_{t}$ within a specified number of resamples. For a given $\delta^{\prime}\in(0,1)$, we set the maximum number of resamples to $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/(1-p))$; then $\mathcal{M}_{t}$ occurs with probability at least $1-\delta^{\prime}/(t+1)^{2}$. Resampling allows the algorithm to further explore the action space to balance regret minimization with accurate reward estimation.
For $a\neq\widetilde{a}_{t}$, i.e., an arm that is not selected in the round $t$, we impute the missing rewards using $\widetilde{\mathbf{x}}_{k}^{\top}\check{\boldsymbol{\mu}}_{t}^{L}$. For $a=\widetilde{a}_{t}$, in contrast, the term $\mathbb{I}(\widetilde{a}_{t}=a)y_{a,t}/\phi_{a,t}$ calibrates the predicted reward to ensure the unbiasedness of the pseudo-rewards for all arms. Given that $\mathbb{E}_{\widetilde{a}_{t}}[\mathbb{I}(\widetilde{a}_{t}=a)]=\mathbb{P}(\widetilde{a}_{t}=a)=\phi_{a,t}$, taking the expectation over $\widetilde{a}_{t}$ on both sides of Eq.˜10 gives $\mathbb{E}_{\widetilde{a}_{t}}[\widetilde{y}_{a,t}]=\mathbb{E}_{t-1}[y_{a,t}]=\widetilde{\mathbf{x}}_{a}^{\top}{{\boldsymbol{\mu}}_{\star}}$ for all $a\in[K]$. Although the estimate $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$ may have a large error, it is multiplied by the mean-zero random variable $(1-\mathbb{I}(\widetilde{a}_{t}=a)/\phi_{a,t})$, which makes the resulting pseudo-rewards defined in Eq.˜10 robust to errors of $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$. The pseudo-rewards can only be computed—and thus can only be used—when the pseudo-action $\widetilde{a}_{t}$ matches the actual action $a_{t}$, that is, when the event $\mathcal{M}_{t}$ occurs. Since $\mathcal{M}_{t}$ occurs with high probability, we are able to compute pseudo-rewards for almost all rounds.
@@
and note that our estimator uses the Gram matrix aggregated over all actions at each round, $\sum_{\tau=1}^{t}\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$, instead of the one built only from the chosen actions, $\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}$. The following theorem guarantees convergence of the estimator, across all arms, to the true parameter, after a sufficient number of exploration rounds.
@@
Let $d_{h}$ denote the dimension of the projected latent rewards defined in Eq.˜6. Then with probability at least $1-3\delta$,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star})|\leq\frac{8\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}},$
for all rounds $t$ such that $t\geq|\mathcal{E}_{t}|$.
Although the DR Lasso estimator leverages $K$-dimensional feature vectors, its error bound depends only logarithmically on $K$. Such fast convergence is typically achieved under classical regularity conditions, including the compatibility condition and the restrictive minimum eigenvalue condition (van de Geer & Bühlmann, 2009; Bühlmann & van de Geer, 2011). Existing Lasso-based bandit approaches (Kim & Paik, 2019; Bastani & Bayati, 2020; Oh et al., 2021; Ariu et al., 2022; Chakraborty et al., 2023; Lee et al., 2025) generally impose these conditions directly on the feature vectors. Our approach, on the other hand, does not require such assumptions, as the augmented features are orthogonal vectors lying in $\mathrm{R}(\mathbf{X})^{\perp}$. Moreover, their average Gram matrix satisfies $\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\geq\min\{1,\lambda_{\min}(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top})\}$. Thus, the convergence rate scales only as $\sqrt{\log K}$ with respect to $K$.
The consistency is proved by bounding the two components of the error in the pseudo-rewards (Eq.˜10): (i) the noise of the reward and (ii) the error of the imputation estimator $\check{\boldsymbol{\mu}}_{t}$. Since (i) is sub-Gaussian, it can be bounded using martingale concentration inequalities. For (ii), the imputation error $\widetilde{\mathbf{x}}_{a}^{\top}(\check{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star})$ is multiplied by the mean-zero random variable $\left(1-{\mathbb{I}(\widetilde{a}_{t}=a)}/{\phi_{a,t}}\right)$ and thus the magnitude of the error can be bounded by $\|\check{\boldsymbol{\mu}}^{L}_{t}-\boldsymbol{\mu}_{\star}\|_{1}/\sqrt{t}$. The detailed proof is deferred to Section˜E.2.
@@
12:   Sample $\widetilde{a}_{t}$ according to  Eq.˜9.
@@
19:   Update $\widehat{\boldsymbol{\mu}}^{L}_{t}$ following Eq.˜11 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}^{L}_{t}$ following Eq.˜8.
@@ 5.1 Robust to Latent Features (RoLF) Algorithm
In this section, we present our algorithm RoLF in Algorithm˜1. In the initialization step, given the observed features, our algorithm constructs a set of orthogonal basis vectors $\{\mathbf{b}_{1}^{\top},\dots,\mathbf{b}_{K-d}^{\top}\}\subseteq\mathrm{R}(\mathbf{X})^{\perp}$ to augment each observed feature vector. The algorithm first selects a candidate action $\widehat{a}_{t}$—either uniformly at random or greedily based on the estimated reward—and then resamples $a_{t}$ and $\widetilde{a}_{t}$ until they match or the maximum number of trials allowed for each round, $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/(1-p))$, is reached. Once the resampling phase ends, the algorithm selects $a_{t}$ and the corresponding reward $y_{a_{t},t}$ is observed. If $a_{t}$ and $\widetilde{a}_{t}$ match within $\rho_{t}$, then both the imputation and the main estimators are updated; otherwise, neither is updated.
Our proposed algorithm does not require knowledge of the dimension of the unobserved features $d_{u}$, nor the dimension of the reward component projected onto $\mathrm{R}(\mathbf{X})^{\perp}$. Although we present the algorithm for fixed feature vectors, the algorithm is also applicable to time-varying feature vectors. In such cases, we estimate the bias caused by unobserved features by augmenting the standard basis. For further details, refer to Appendix˜D.
@@ 5.2 Regret Analysis
Let $d_{h}$ denote the number of nonzero coefficients in the representation of the projected latent reward as defined in Eq.˜6. Then for any $\delta\in(0,1)$ and $p\in(1/2,1)$, with probability at least $1-6\delta$, the cumulative regret of RoLF is bounded by
$\begin{split}\mathrm{Reg}(T)&\leq 2\cdot 8^{3}K^{3}(1-p)^{-2}\log\frac{2KT^{2}}{\delta}\\ &\quad+\frac{4\sqrt{T}}{K-1}+2\sqrt{2T\log\frac{2}{\delta}}+4\delta\\ &\quad+\frac{32\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}}.\end{split}$
To the best of our knowledge, Theorem˜3 provides the first regret bound that converges faster than $\widetilde{O}(\sqrt{KT})$, specifically for algorithms that account for unobserved features without relying on any structural assumptions. Assuming $\|\widetilde{\mathbf{x}}_{a}\|_{\infty}\leq 1$ (instead of $\|\widetilde{\mathbf{x}}_{a}\|_{2}\leq 1$), $\widetilde{\sigma}^{2}_{\min}$ and $\widetilde{\sigma}_{\max}^{2}$ are constant independent of $d$ or $K$. Note that the number of rounds required for the exploration phase is $O(K^{3}\log KT)$, which grows only logarithmically with the time horizon $T$. The factor $K^{3}$ is irreducible, as the algorithm needs information about all $K$ biases from the missing features. By employing the Gram matrix of the augmented feature vectors over the action space $\mathcal{A}$, $\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$, when combined with DR estimation, we reduce the required exploration phase time from $O(K^{4}\log KT)$ to $O(K^{3}\log KT)$, thereby improving the sample complexity by a factor of $K$.
As demonstrated in Theorem˜2, the last term on the right-hand side of the regret bound is proportional to $\sqrt{d+d_{h}}$ rather than $\sqrt{K}$, resulting in an overall regret bound of $O(\sqrt{(d+d_{h})T\log KT})$. The specific value of $d_{h}$ is determined by the relationship between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$, as discussed in Section˜4.1. Table˜2 summarizes how the regret bound varies under different relationships between these subspaces. The formal proof of Theorem˜3 is deferred to Section˜E.3.
@@ 6.1 Experimental Setup
In this experiment, we simulate and compare our algorithms, Algorithm˜1 and Algorithm˜2 (Appendix˜C), with baseline OLB algorithms: LinUCB (Li et al., 2010) and LinTS (Agrawal & Goyal, 2013). These algorithms adopt UCB and Thompson sampling strategies, respectively, assuming a linear reward model based on observed features. We also include DRLasso (Kim & Paik, 2019) since our algorithm employs DR estimation with a Lasso estimator, and UCB($\delta$) (Lattimore & Szepesvári, 2020), an MAB algorithm that ignores features, to assess the effectiveness of using feature information, even under partial observability.
@@
In the simulation, after constructing the true features $\mathbf{z}_{a}\in\mathbb{R}^{d_{z}}$ and observed features $\mathbf{x}_{a}\in\mathbb{R}^{d}$, we apply singular value decomposition (SVD) to the observed feature matrix $\mathbf{X}$ to derive orthogonal basis vectors $\{\mathbf{b}_{1}^{\top},\dots,\mathbf{b}_{K-d}^{\top}\}$ orthogonal to $\mathrm{R}(\mathbf{X})$. These vectors are then concatenated with $\mathbf{X}$ to form the augmented feature matrix. The reward parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$ is sampled from $\mathrm{Unif}(-1/2,1/2)$, and the rewards are generated following Eq.˜2. For the hyperparamters in our algorithms, the coupling probability $p$ and the confidence parameter $\delta$, are set to $0.6$ and $10^{-4}$, respectively. The total decision horizon is $T=1200$. Throughout the experiments, we fix the number of arms at $K=30$ and the dimension of the true features at $d_{z}=35$, ensuring $d_{z}\geq d$ to cover both scenarios. Further details on the experimental setup are provided in Appendix˜B.
@@ 6.2.1 Scenario 1
In this scenario, $d$ is set to $\lfloor d_{z}/2\rfloor$, so that the agent observes only about half of the full feature dimension. From Figure˜2, we observe that the baseline OLB algorithms—LinUCB, LinTS, and DRLasso—perform worse than our algorithms in terms of both the level and robustness of cumulative regret. This pattern is consistently observed across all three cases, implying that linear bandit algorithms fail to identify the optimal arm in the presence of unobserved components, as they cannot capture the portion of the reward associated with the unobserved features. However, our algorithms show almost the same performance regardless of the relationship between $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$, indicating the robustness to variations in feature observability structure.
Note that for all cases our algorithms—RoLF-Lasso (Algorithm˜1) and RoLF-Ridge (Algorithm˜2)—exhibit a sharp decline in the growth rate of cumulative regret after a certain number of rounds. This behavior is primarily due to the forced-exploration phase built into the algorithms, which ensures sufficient coverage of the action space in the early stages. Following this phase, the resampling and coupling strategies further enhance the efficiency of the DR estimation, leading to slower regret accumulation over time. This pattern is commonly shown from experimental results of other bandit algorithms employing the forced-exploration strategy (Goldenshluger & Zeevi, 2013; Hao et al., 2020; Chakraborty et al., 2023).
Furthermore, in Case 2 (Figure˜2(b)), the cumulative regret grows more slowly than in other cases. This behavior is explained by the relationship $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$, implying the reward components projected onto $\mathrm{R}(\mathbf{X})^{\perp}$ can be fully expressed by the observed features. As a result, the augmented features fully capture the underlying reward structure, enabling faster convergence relative to the other scenarios.